OpenStudy (anonymous):

Evaluate 7 − 2y, if y = 5

OpenStudy (theviper):

-3

OpenStudy (anonymous):

7-2(5)=-3

OpenStudy (theviper):

\[7-2(5)\]=7-10=-3